OpenStudy (anonymous):

calculus help.. can someone tell me if my answer is correct??? find the derivative of the function. y=x^2e^-x

OpenStudy (anonymous):

y(prime)=2x+e^-x*-1 my answer=2x-e^-x u=-x u(prime)=-1

OpenStudy (anonymous):

\[x^2 e^{-x} = \frac{x^2}{e^x}\] Now, do the quotient rule.

OpenStudy (anonymous):

oooooooooooooooooooooooooooooooooooo okay that makes sense...so my answer is wrong hugh?

OpenStudy (anonymous):

oh hell no quotient rule is a pain

OpenStudy (anonymous):

Yeah, it seems so.

OpenStudy (anonymous):

lololololololo

OpenStudy (anonymous):

can i use the product rule instead?

OpenStudy (anonymous):

\((fg)'=f'g+g'f\) just like yesterday

OpenStudy (anonymous):

in this case \(f(x)=x^2,f'(x)=2x,g(x)=e^{-x}, g'(x)=-e^{-x}\)

OpenStudy (anonymous):

plug and chug

OpenStudy (anonymous):

awww u remembered! hahah okay i got it...but one question....my answer was way off right?

OpenStudy (anonymous):

way off, right

OpenStudy (anonymous):

:{.................okay :]

OpenStudy (anonymous):

so my answer is this right? y(prime)=2x(e^x)-e^-x (x^2)

OpenStudy (anonymous):

Quotient rule wasn't so hard. Just saying. \[\frac{d}{dx} \frac{f(x)}{g(x)} = \frac{f'(x)g(x) - f(x)g'(x)}{(g(x))^2}\] It looks complicated but all you have to do is derivative factors then plug in, just like the way satellite did.

OpenStudy (anonymous):

functions*

OpenStudy (anonymous):

lo-dee-hi-hidee-lo-all-over-lo^2

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so my answer is right, right?

OpenStudy (anonymous):

Well, my answer is \[y \prime = \frac{2x - x^2}{e^x}\]

OpenStudy (anonymous):

same thing.....thanks!

OpenStudy (anonymous):

Yeah your answer is right, but you can simplify it.

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

No problem!